OpenStudy (anonymous):

find limit as x approaches negative infinity x / (x-sqrt. (x^2 +7))

myininaya (myininaya):

\[\lim_{x \rightarrow -\infty}\frac{x}{x-\sqrt{x^2+7}}\]

myininaya (myininaya):

So do you know what to divide both top and bottom by?

OpenStudy (anonymous):

oh yeah! i get it now! thank you :)

myininaya (myininaya):

don't forget \[\sqrt{x^2}=|x|=-x \text{ since } x<0\]